Circular No. 7526

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 W1 (UTSUNOMIYA-JONES) On Nov. 19, S. Nakano, Sumoto, Japan, reported the visual discovery

on Nov. 18 by Syogo Utsunomiya (Aso, Kumamoto; 25×150 binoculars) of a possible comet with coma diameter 5′ moving rapidly southeastward in Vela. Attempts by several observers (including A. Hale, D. Seargent, J. Biggs, T. Urata, and J. Kobayashi) to confirm the object, at the request of Nakano and the Central Bureau, were unsuccessful. On Nov. 25, A. C. Gilmore (Mount John University Observatory) reported the visual discovery of an apparent comet by Albert F. Jones (Nelson, New Zealand, 0.078-m f/8 refractor, 30×) while observing the variable star T Aps at dawn; Jones reported the comet as being diffuse with coma diameter ∼ 4′ in morning twilight. The possibility that Jones’ object might be the same as that reported by Utsunomiya was explored by the Central Bureau, and a search ephemeris from plausible parabolic orbital elements fitted to the Nov. 18 and 25 approximate positions was circulated to numerous southernhemisphere observers. Confirming CCD astrometry was made by Gilmore with the 1.0-m f/7.7 reflector at Mt. John.

2000 UT  α_{2000}  δ_{2000}  m_{1} Observer
Nov. 18.817  9^{h}04.^{m}0  −41^{o}22′  8.5 Utsunomiya
18.833  9 04.2  −41 27  8.5  _{”}
18.847  9 04.4  −41 31  8.5  _{”}
25.64  13 49.8  −77 05  8  Jones
26.42295 15 36 22.^{s}31 −76 23 51′′.5  Gilmore
26.42506 15 36 37.51 −76 23 31.7  _{”}
26.42772 15 36 57.31 −76 23 05.1  _{”}
26.42998 15 37 13.94 −76 22 43.8  _{”}
26.55906 15 52 38.76 −76 00 58.0  _{”}
26.56010 15 52 46.20 −76 00 47.4  _{”}
26.56126 15 52 54.24 −76 00 34.8  _{”}
26.56207 15 52 59.84 −76 00 25.6  _{”}

The following preliminary parabolic orbital elements by B. G. Marsden are heavily weighted toward the Mt. John accurate positions:

T = 2000 Dec. 26.56 TT  ω = 51.^{o}52
Ω = 10.81 }2000.0
q = 0.3214 AU  i = 160.16
2000 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green